OpenStudy (anonymous):

Use elimination to solve the system. What is the value of y? 3x + 5y = 16 and 8x - 5y =28

jimthompson5910 (jim_thompson5910):

Add the two equations to get 3x + 5y = 16 8x - 5y =28 ------------ 11x+0y = 44 So 11x = 44 Now solve for x 11x = 44 11x/11 = 44/11 x = 4 From here, use this x value in either equation to find y. I'm going to pick on the first equation 3x + 5y = 16 3(4) + 5y = 16 ... Plug in x = 4 12 + 5y = 16 5y = 16 - 12 5y = 4 y = 4/5 So the value of y is 4/5

jimthompson5910 (jim_thompson5910):

Hopefully that gives you a pretty good idea of how elimination works so you can use it to solve other similar problems.

OpenStudy (anonymous):

but my options are: 4 -6 0.8 5

jimthompson5910 (jim_thompson5910):

Using a calculator, or by performing long division, \[\Large \frac{4}{5} = 0.8\]

jimthompson5910 (jim_thompson5910):

So y = 0.8 as well (along with y = 4/5)

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw